OpenStudy (bernalcheryl):

i want to know if this problem is correctly solved if not please help me . Use the distributed law to factor 14u+49v 7.2 7.7 7(2u+7v)

OpenStudy (whpalmer4):

Testing your factoring is easy: just multiply it out! \[7(2u+7v) = 7*2u + 7*7v = 14u + 49v\]

OpenStudy (bernalcheryl):

so it is correct becuase i kinda did same thing you just show me